OpenStudy (anonymous):

the ratio of the side lengths of a quadrilateral is 4: 3 : 8 : 5. Its perimeter is 120 centimeters. What is the length of the longest side? 6cm 30cm 24cm 48cm

OpenStudy (amistre64):

n(4+3+8+5) = 120

OpenStudy (amistre64):

n(20) = 120 n = 120/20 = 6 6(8) = 48

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (amistre64):

youre welcome :)